Mauro Products distributes a single product, a woven basket whose selling price is $10 per unit and whose variable expense is $8 per unit. The company’s monthly fixed expense is $6,000. Required:

1. Calculate the company’s break-even point in unit sales.

2. Calculate the company’s break-even point in dollar sales. (Do not round intermediate calculations.)

3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do not round intermediate calculations.)
Answer #1

1) Break even point = Fixed cost/Contribution margin per unit = 6000/(10-8) = 3000 Units

2) Break even sales = 3000*10 = $30000

3) New break even point = 6600/2 = 3300 Units

Break even sales = 3300*10 = $33000